A blue spruce grows an average of 6 in per year. A hemlock grows an average of 4 in per year. If a blue spurce is 4 feet tall and a hemlock is 6 feet tall, when would you expect the trees to be the same height?
Help, i hate written problems.

Answer by checkley75(3666)

(48+6X)=(72+4X)
48+6X=72+4X
6X-4X=72-48
2X=24
X=24/2
X=12 YEARS THE TWO TREES WILL BE @ THE SAME HEIGHT.